OpenStudy (anonymous):

The function f(t) = 7 cos(pi over 4t) + 12 represents the tide in Light Sea. It has a maximum of 19 feet when time (t) is 0 and a minimum of 5 feet. The sea repeats this cycle every 8 hours. After two hours, how high is the tide? @ganeshie8

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

tough question. idk.

ganeshie8 (ganeshie8):

plugin \(t = 2\)

ganeshie8 (ganeshie8):

\(f(t) = 7 \cos(\frac{\pi}{4}t) + 12 \)

ganeshie8 (ganeshie8):

\(f(2) = 7 \cos(\frac{\pi}{4}*2) + 12 \) \(f(2) = 7 \cos(\frac{\pi}{2}) + 12 \) \(f(2) = 7 *0 + 12 \) \(f(2) = 12 \)

OpenStudy (ready):

help!

OpenStudy (solomonzelman):

I would suggest to open a new question. If you have one question open you can close it and open a new one.